# RealitySummary: On-Demand Mixed Reality Document Enhancement using Large Language Models

### v1

@@ 5.1. Method
We recruited participants (P1-P12, 6 male, 6 female; Ages = 19-35 years) through word-of-mouth and they were compensated with \$15 for their one-hour participation in the study. The study was conducted with Microsoft Hololens 2 in a lab environment. All the participants had verbal and written fluency in English.
Before starting the study, the authors familiarized participants with AR and the Hololens 2, then introduced RealitySummary and each of the document enhancements described in the previous section. After this introduction, we provided three reading tasks using: 1) an A4 size magazine in physical format, 2) a digital document of the participant’s choice, and 3) an A3 size book. Participants’ digital documents included 3 books, 2 blog posts, 2-course assignment PDFs, 2 Wikipedia articles, 2 magazine articles, and 1 news article which they viewed on either a PC (4/12), tablet (5/12), or laptop (3/12). We asked participants to use RealitySummaryto aid with their reading, using whatever combination of enhancements best suited the current document. Throughout this 30-minute period, we encouraged participants to think aloud, documenting their experiences with the system. After the tasks, we conducted semi-structured interviews and asked them to complete a Likert scale questionnaire, including a System Usability Scale (SUS) protocol (Brooke, 1995). Together, the questionnaire and interview took roughly 20 minutes.
Afterward, we conducted a thematic analysis by first transcribing the think-aloud sessions and interviews, then individually familiarizing ourselves with the content and organizing data into rough categories. Two authors then individually searched for broader groups and patterns and clustered these into themes. The writing team then iterated to further refine the identified themes.
@@ 5.2. Results
Participants responded positively to RealitySummary (Figure 5), with most participants either agreeing or strongly agreeing that the the prototype was easy to use (10/12), easy to understand (11/12), and well-integrated with the task (10/12). Similarly, nearly all participants reported that the summary approaches provided by RealitySummary helped with comprehension (12/12), were relevant to the document content (11/12), improved comprehension speed (10/12), and augmented their reading experience (10/12). Finally, the System Usability Scale (SUS) returned an overall composite score of 71, suggesting a reasonable level of usability for an early-stage prototype.
Figure 5. Usability study questionnaire responses.
5.3. Findings
@@
Our analysis identified five themes related to participants’ experience comprehending and navigating documents, as well as their sense of focus and control. We also discuss our participants’ comparisons of the visual and text-based summarization techniques and layouts implemented in RealitySummary.
Comprehension
@@
Document enhancements can improve readers’ sense of comprehension, and readers value the ability to see multiple types of enhancements simultaneously. All the users in the study felt that the system improved their reading comprehension. Out of the 12 participants surveyed, 11 reported that they were able to understand the document without actually reading it, solely by relying on the contextual information provided by RealitySummary. P9 even stated, “I didn’t even read the document but I understand what it is about.” Furthermore, multiple participants (P7, P8, P11, P12) argued that the combination of multiple features like summaries, timelines, keywords, images, and people cards provided them with a more comprehensive understanding of the document, allowing them to engage with it from different angles. This allowed them to gain multiple perspectives on the same topic, leading to a deeper comprehension of the material. All participants mentioned that RealitySummary helped them comprehend documents more quickly than their typical approaches (like visually scanning a physical book and manually searching for a related summary or timeline on a phone/PC). For example, after perusing the first magazine, P12 noted that “I’ve not read this magazine thoroughly but, it seems that these features can significantly help me understand what this document is about in less amount of time.”.
Navigation
@@
Both structured overviews and in-context highlights provide useful entry points that can make it easier to navigate documents in a top-down fashion, even within individual pages. Broadly, participants spoke positively about how RealitySummary helped them navigate documents. Four participants (P3, P4, P6, P8) specifically reported using structured navigation elements like timelines to get a better understanding and structural flow of the whole text, with P4 noting that “the timeline helped me organize and navigate my thoughts very quickly while reading the text” (P4). Meanwhile, at least a third of participants (P5, P6, P9, P10) also mentioned using lower-level augmentations to help navigate the document— highlighting how the proximity of annotations and content allowed them to use keyword highlights and information cards as visual anchors. Additionally, several participants found the connection lines between highlighted words and their corresponding features to be a valuable tool for navigating the document—with P8 stating, “the words in the documents are highlighted […] and connected with all the features […] it is very helpful. It is also pretty helpful that I can see the word ‘Yosemite’ and where it is on the map around the text!”.
In fact, ten different participants reported that the system allowed them to focus their reading by taking a “top-down” approach—first examining high-level ideas and terms surfaced by RealitySummary, then gradually transitioning into reading lower-level content where necessary. P8 highlighted that “[the enhancements] give you a foundation before you actually jump inside [the document]”. This allowed participants more leeway to prioritize their reading to “focus on things that you’re more unsure about rather than just kind of trying to find external information” (P1).
Interestingly, several participants also reflected positively on the fact that summaries and keywords surfaced in our prototype were drawn only from the currently visible page, rather than the document as a whole. P5 reported that, “I like page by page summary—having a summary for each page does help make it feel less dense compared to a summary of the whole document”. Similarly P6 stressed that, “I like that it (system) gives you keywords for each page and not as a whole.
Visual vs. Text Enhancements
Both visual and text-based enhancements can benefit readers, with visuals providing supporting high-level overviews and providing context, and text summaries capturing more complex concepts. While our participants responded positively to all of the document enhancements in the prototype, they reported that they preferred text features when trying to gain a deeper understanding of a particular topic or content within the document. Conversely, participants generally felt that visual features improved their comprehension speed, increased engagement, and improved memorability.
@@
All 12 participants ranked the visual summarization techniques—including image cards, people bios, and maps—as helpful, with eight participants specifically highlighting the value of images and bio cards. Many participants (P1, P4, P6, P7, P9, P10, and P11) agreed that the visual augmentations helped them understand the document faster compared to just reading the text, noting that, “seeing the visuals helped me grasp what’s happening in the document right away” (P11). Several participants (P7, P8) also stressed the potential for visuals to increase their engagement with the text, with P7 stating, “I’m not a big fan of reading so having visuals are helping me assist in reading this”. These participants also highlighted the potential for visuals to increase memorability with P8 noting that for historical documents the people cards could “help them put a face to a name”. Similarly, P7 noted that while text-based summaries might contribute to greater understanding, “graphical stuff makes me remember it for a longer time”.
@@
Placing enhancements in the space around the document using mixed reality can successfully reduce the potential for visual clutter and distraction. Because they can sit alongside and even on top of document content, summaries, information cards, and other document enhancement have the potential to occlude or distract attention from the document. However, reducing the number of enhancements or moving them further from the text may limit their utility. In our case, several participants (P7, P10, P11) specifically highlighted how our around-document placement of the layouts integrated nicely with the documents themselves—calling them natural and intuitive to use. P10 reported that “[the system] is like a natural extension of the book. It is very easy to follow through.”. Similarly, four participants (P1, P2, P5, P11) praised the fact that the augmentations did not disrupt their visual focus on the document, with P5 noting that “it is not cutting into your field of view and it’s not invasive at all”. Moreover, 10 participants also remarked positively on the connecting lines linking document text to cards around around the document—indicating that they were easy to follow but did not impede readability.
6. In-the-Wild User Study
@@
To explore additional real-world use cases, we also conducted an in-the-wild study with eleven participants. In contrast to the previous usability study, which focuses on standardized reading tasks in a lab setting, this study instead focused on examining the versatility and applicability of our system across uncontrolled environments and diverse reading materials. Our goal was to uncover unique advantages and potential use cases of employing mixed reality and AI-generated assistants in everyday settings. For this study, the system was deployed on an Apple Vision Pro, which provides considerably higher rendering quality than the Hololens 2. To accommodate diverse applications and reduce the obtrusiveness of additional content in everyday scenarios, the user interface was refined, limiting augmented content to a text box and an image slot.
Figure 6. In-the-wild study.
@@ 6.1. Method
We recruited 11 participants (W1-W11, 7 male and 4 female; Ages 19-32 years) from the local community. They were compensated with \$25 for their participation in a 1.5-hour study, which consisted of a one-hour experiment and a half-an-hour interview. The study was conducted in various locations at a local university, including a library, cafeteria, store, hallway, classroom, and multiple lab settings.
Before the experiments, we collected participants’ demographics and their reading habits, which were input into the system as the prior knowledge of ChatGPT. Participants familiarized themselves with Apple Vision Pro by doing the eye tracking calibration setup. Unlike structured tasks or prepared reading materials, participants were encouraged to utilize our system with any text content they encountered during the experiments, including documents, books, papers, posters, online news on mobile devices, and other materials in their environment. Throughout this one-hour experiment, we also encouraged participants to think aloud about their experiences. After the experiment session, we conducted semi-structured interviews, followed by a Likert scale questionnaire similar to the usability study. Both the interview and questionnaire took approximately 30 minutes. As in the earlier usability study, we then analyzed participants’ questionnaire responses and reflections to identify high-level trends and opportunities for augmented reading tools.
@@ 6.2. Results
We gathered feedback from participants on various aspects including the reading experience (immersiveness, understanding, distraction), perceptions of the system (content and query comprehension, trustworthiness, and privacy concerns), and overall usability (content placement and comfort using the MR headset). In general, most participant either agree or strongly agree that RealitySummary deepened their understanding of the reading content (7/11), RealitySummary demonstrated a deep understanding of the content they were reading (7/11). Participants reported feeling more immersed and interested while reading, with an average rating of 3.45 (SD = 1.04), and they perceived the system as effectively interpreting their queries (M = 3.27, SD = 1). Besides, their responses on the content placement within the system were also overall positive (M = 3.36, SD = 1.29).
Participants had varied opinions on whether the system became distracting throughout the reading process (M = 2.55, SD = 1.37), and whether they tended to trust the accuracy and reliability of contents provided by the system (M = 2.82, SD = 1.25). Privacy concerns were also noted by the majority (6/11), and discomfort using the headset was prevalent among the respondents (8/11). These findings are detailed in Figure 7.
@@
6.3. Findings
@@
Benefits of Always-On Assistant
The exploration of such diverse reading materials can be attributed to the always-on capability of the mixed reality reading assistant. A key advantage of this always-on feature is its support for implicit interactions, offering a distinct user experience compared to ChatGPT on smartphones or laptops, which require explicit effort to engage. This implicit and automatic content capture greatly expands the range of applications, beyond what participants initially considered possible. The always-on capturing feature eliminated the need for manually supplying context, saving considerable time. For example, W9 experienced a smoother workflow when working on her business course project with RealitySummary compared to conventional screen-based methods. She was able to concentrate on her project, without having to switch between smartphones or laptops to familiarize ChatGPT with the context.
@@
Beyond the always-on implicit assistance, participants appreciated the unique affordances offered by mixed reality interfaces, such as the opportunities for spatial and tangible exploration. In a library setting, for example, participants could naturally navigate through bookshelves (spatial exploration), physically pick up a book (tangible interaction), and browse through its pages to automatically receive summaries (always-on assistance). W5 and W10 even leveraged our system to provide summaries or recommendations for entire collections of books on a bookshelf, not just content from a single page. Furthermore, W1 demonstrated the capability to synthesize information across multiple books on a table, requesting summaries or connections between various books and pages. Such observations highlight how mixed reality interfaces enable spatial and tangible interactions that go beyond the limitations of traditional screen-based interfaces, which can only capture and summarize visible content.
@@
We observed several unique and unexpected use cases that took advantage of the LLM’s capacity for free-form questioning and answering. For instance, when reading a physics textbook, W8 asked the system to generate ten quizzes to help him understand quantum mechanics, transforming the textbook into a more active and personalized learning tool. Other common usages included translations (W4, W5, W6), learning aids (W4, W8), and personalized content recommendations (W5, W10). For example, applied the system to his handwritten calculus notes to find answers to particular questions. Beyond traditional reading applications, the AI’s capabilities supported practical decision-making; W4 compared nutritional information of products on store shelves and menu items to determine healthier options, while W5 used it for guidance on waste disposal and printer operations. The LLM’s ability to process a broad spectrum of inquiries significantly expanded the system’s applicability and utility.
@@
Our system offers both 1) proactive assistance through automatic summary generation and 2) on-demand assistance where users manually ask questions. Participants shared their insights on these two functionalities by comparing the benefits and limitations. Proactive summaries were appreciated for providing a useful initial overview of unfamiliar topics. Preferences for proactive summaries varied with reading habits. For example, infrequent readers like W6 and W10 found these proactive summaries helpful in deciding whether to read the content or understand it without reading the full text. On the other hand, W10, who is a more frequent reader, found the constant summary updates unnecessary. In contrast, on-demand question answering allowed for a more focused engagement on specific topics. For instance, W1 enjoyed the ”rapid-fire” questioning with a physics textbook, asking more questions than with traditional ChatGPT saying ”You have to describe it properly you have to specifically say what you’re asking about what you’re reading, so that kind of all of this abstracted in this system is like frees up your mind to just focus on the curiosity part Like rapid-fire questions and get rapid-fire answers”. W2 valued how questioning was woven into reading, preserving thought flow without the disruption of toggling between reading and ChatGPT interfaces. Overall, participants favored the on-demand feature due to its applicability to diverse situations, yet many appreciated having both options for different purposes: proactive features for implicit skimming and on-demand for addressing more specific questions.
Trust of AI-Generated Content
Many participants approached the system’s responses with caution, often verifying the answers with the reading material. They highlighted the necessity of consulting additional sources for in-depth or technical readings when depending on the system. There was a general agreement among participants about their reluctance to fully trust an always-on reading assistant in their daily activities. Conversely, we noted some intriguing discussions regarding how participants perceive our system, either as a mere tool or more like a personal companion. While most participants engaged with the system as a conventional chatbot, some regarded it more as a personal companion than a mere tool. Notably, W3 and W5 described the system as a friend for discussing and inquiring about their readings. The system’s ability to provide always-on immediate responses facilitates this personification.

### v2

@@ 4.2. Method
The usability study involved 12 participants (P1-P12, 6 male and 6 female, aged 19-35) recruited via word-of-mouth and compensated \$15 for their one-hour participation. The study aimed to evaluate the system’s usability, explore participants’ preferences for different document enhancement strategies, and identify limitations for future iterations with real-world use in mind. Conducted in a lab environment with Microsoft HoloLens 2, participants were first introduced to AR and the HoloLens 2, followed by an overview of the system and its document enhancements. They completed three reading tasks using different document formats (physical, digital, and a book) and utilized the system’s enhancements as they preferred. The participants’ digital documents included books, blog posts, course assignments, Wikipedia articles, magazine articles, and news articles viewed on various devices.
Throughout the study, participants were encouraged to think aloud, and their experiences were documented. Following the tasks, semi-structured interviews and a Likert scale questionnaire, including a System Usability Scale (SUS), were conducted. Thematic analysis was performed on the transcriptions of the think-aloud sessions and interviews, resulting in the identification and refinement of key themes.
@@ 4.3. Results
Overall, participants responded positively to RealitySummary, with most participants agreeing that the prototype was easy to use (M=4.3/5, SD=0.98), easy to understand (M=4.5/5, SD=0.67), and well-integrated with the task (M=4.3/5, SD=0.77) with 5-point Likert-scale questionnaire. Similarly, nearly all participants reported that the summary approaches provided by RealitySummary helped with comprehension (M=4.6/5, SD=0.514), were relevant to the document content (M=4.7/5, SD=0.65), improved comprehension speed (M=4.4/5, SD=0.79), and augmented their reading experience (M=4.7/5, SD=0.65). Finally, the System Usability Scale (SUS) returned an overall composite score of 71, suggesting a reasonable level of usability for an early-stage prototype.
4.4. Insights
@@
Our analysis identified several themes related to participants’ experience in comprehending and navigating documents. We also explore participants’ comparisons of the visual and text-based summarization techniques as well as preferences for various augmentation techniques.
Enable Multi-Angle Content Understanding for the Current Page
@@
RealitySummary’s document enhancements can significantly improve readers’ comprehension. Several participants (P7, P8, P11, P12) highlighted that features such as summaries, timelines, keywords, images, and people cards provided a more comprehensive understanding of the document, allowing them to engage with the content from multiple angles and perspectives. All participants mentioned that RealitySummary enabled them to grasp content more quickly than with their usual methods, such as visually scanning a physical book or manually searching for timelines on a phone or PC. For instance, after reviewing the first magazine, P12 remarked, I haven’t read this magazine thoroughly, but it seems these features can significantly help me understand what it’s about in less time.” Interestingly, several participants appreciated that the summaries and keywords were drawn only from the currently visible page rather than from the entire document. P5 reported, I like the page-by-page summary—having a summary for each page does help make it feel less dense compared to a summary of the whole document.” Similarly, P6 emphasized, “I like that it (the system) gives you keywords for each page and not for the whole document.
Facilitate Efficient, Top-Down Navigation
@@
Both structured overviews and in-context highlights provide useful entry points that make it easier to navigate documents from a top-down perspective, even within individual pages. Broadly, participants spoke positively about how RealitySummary helped them navigate documents. Four participants (P3, P4, P6, P8) specifically mentioned using structured navigation elements like timelines to better understand the structure and flow of the text, with P4 noting that the timeline helped me organize and navigate my thoughts very quickly while reading the text.” In fact, ten participants reported that the system allowed them to focus their reading by taking a “top-down” approach—first examining the high-level ideas and terms, then transitioning into reading the lower-level content as needed. P8 highlighted, “[the enhancements] give you a foundation before you actually jump inside [the document].” This approach enabled participants to prioritize their reading, allowing them to “focus on things that you’re more unsure about” (P1).
Support Quick Understanding with Visual Aids
@@
Both visual and text-based enhancements offer distinct benefits to readers—visual aids provide quick, high-level overviews, while text summaries capture more complex details. Some participants preferred text-based features for gaining a deeper understanding of specific topics, while others found that visual elements improved their comprehension speed, engagement, and retention. All 12 participants rated the visual summarization techniques as helpful, with eight specifically highlighting the usefulness of images and bio cards. Several participants (P1, P4, P6, P7, P9, P10, P11) noted that visual augmentations allowed them to grasp the document’s content more quickly compared to reading text alone, with P11 stating, “Seeing the visuals helped me grasp what’s happening in the document right away.” Participants (P7, P8) also emphasized that visuals enhanced their engagement with the material, with P7 commenting, I’m not a big fan of reading, so having visuals helps me engage with the content.” Additionally, participants highlighted the potential of visuals to improve memorability, with P8 noting that for historical documents, people cards helped put a face to a name.” Similarly, P7 observed that while text-based summaries aided understanding, “graphical elements help me remember it for a longer time.”
4.5. Limitations
Slow and Unreliable Document Tracking
Document tracking enables unique features like keyword highlighting and around-document placement. However, the primary limitation of the current prototype is the instability of its document tracking. Overall, participants reported dissatisfaction with the system’s slow, inaccurate, and unreliable tracking capabilities. Unlike the OCR process, document tracking involves a heavier and slower workflow, requiring rectangle detection, image extraction, and image upload for Vuforia processing. Even with a perfect image target, tracking performance varies based on several factors, such as lighting conditions, hand occlusion, and viewing angles. Furthermore, while the tracking performed better for visually rich documents, it struggled with text-heavy documents, as it does not have any distance features, making consistent tracking difficult over time. Even when tracking functioned properly, participants frequently experienced disruptive flickering, which negatively affected the user experience. Additionally, the fixed placement of augmented information on documents posed a challenge for long-term use, as participants expressed the need for more flexibility in adjusting the location of the augmentations.
@@ 4.5. Limitations
Several participants (P7, P10, P11) noted that around-document augmentation helped them focus on their reading but emphasized that the number of displayed cards and their placements should be adjusted based on the user’s needs and context. Displaying too many cards at once can lead to visual clutter and divert attention from the document itself. We observed that participants might become overly reliant on the generated content rather than engaging directly with the reading material. Whether positive or negative, 11 out of 12 participants reported that they could understand the document without actually reading it, relying entirely on the contextual information provided by RealitySummary. P9 even stated, “I didn’t even read the document, but I understand what it’s about.” Currently, the system automatically generates various types of cards when relevant, but allowing users to reduce the number of visible cards based on their preferences could help manage visual clutter and maintain focus.
Fixed Prompt and Augmentation Types
Another key issue was the inability of participants to query the generated summaries. The fixed prompt approach used for generating the different types of cards limited the system’s flexibility in addressing broader use cases and contexts. Several participants (P1, P3, P5) expressed frustration with this constraint, as they had specific questions they wanted to ask but were restricted to the preset summary types provided. While these different types of augmentations offered different benefits, those advantages were only realized when the generated content was relevant and accurate. Participants also noted that specific aids, like timelines, comparison tables, and keyword lists, were not always helpful or contextually appropriate. On the other hand, participants generally appreciated the broader usefulness of the general summaries and visual information cards, which were more adaptable to varying contexts.
Applicability Beyond Controlled Environments
Finally, the controlled nature of the study limits its ability to assess the broader applicability and real-world uses. While the initial findings are encouraging, they do not fully reflect how the system might perform in diverse, unstructured environments. As a result, transitioning to in-the-wild studies will be essential to understanding how document augmentations work in everyday scenarios and to gain insights from prolonged, regular use in real-world settings.
In conclusion, a more refined version of the system, featuring flexible augmentation placement and query-able summaries, is essential for better evaluating its effectiveness in real-world scenarios.
5. Study 2 RS-WILD: In-the-Wild Study with the Second Prototype
Figure 6. Examples from our in-the-wild study.
@@ 5. Study 2 RS-WILD: In-the-Wild Study with the Second Prototype
To explore real-world use cases, we conducted an in-the-wild study with eleven participants. Unlike the previous usability study, which focused on standardized reading tasks in a controlled lab setting, this study aimed to examine the versatility and applicability of our system in uncontrolled environments and with diverse reading materials. The goal was to uncover unique advantages and potential use cases for using mixed reality and AI-generated assistants in everyday settings. For this study, the system was deployed on an Apple Vision Pro, offering considerably higher rendering quality and a broader field of view compared to the Hololens 2. This contributes to better readability, particularly for text summarization, by using a less distractive transparent background. Based on the user feedback from the usability study, the user interface was simplified with a single text box and an image slot, which could accommodate any response type given the user’s custom query. This feature would better suit diverse applications and minimize the intrusiveness of additional content in daily use.
5.1. Implementation
The system design for the Vision Pro follows a similar approach to the Hololens but with key adjustments. Since the Vision Pro does not yet allow direct access to the camera view, we implemented a workaround for text extraction. First, we screencast the Vision Pro to an external iPad, which is then connected to a MacBook Air to stream the feed via QuickTime Player’s mirroring feature. A web-based interface on the Mac captures the live stream from the Vision Pro. Using the captured feed, we employed Google OCR to extract text, which was then processed by GPT-4 via the OpenAI API. The Vision Pro application, developed in Swift, featured a single movable card with an image/map slot. After querying the LLM, data was transmitted from the web-based interfaces to Vision Pro via Firebase’s real-time database.
Additionally, a separate web interface with speech recognition was run on the user’s smartphone, allowing them to query the system via voice (using Web Speech Recognition) and ask follow-up questions. This enabled users to request specific summary types or submit open-ended queries about the text they were reading. The web-based interfaces, built with React.js and hosted on Glitch, were accessed by participants through their phones.
@@ 5.2. Method
We recruited 11 participants (W1-W11, 7 male and 4 female; ages 19-32 years) from the local community. They were compensated \$25 for participating in a 1.5-hour study, which included a one-hour experiment and a 30-minute interview. The study took place in various locations at a local university, such as a library, cafeteria, store, hallway, classroom, and multiple labs.
Before the experiment, we collected participants’ demographic information and reading habits, which were input into the system as prior knowledge for ChatGPT. Participants familiarized themselves with the Apple Vision Pro by completing the eye-tracking calibration setup. Unlike structured tasks or prepared reading materials, participants were encouraged to use the system with any text they encountered during the study, including documents, books, papers, posters, online news on mobile devices, and other materials in their environment. Throughout the one-hour experiment, participants were encouraged to think aloud about their experiences. After the experiment, we conducted semi-structured interviews, followed by a Likert scale questionnaire similar to the one used in the earlier usability study. Both the interview and questionnaire took approximately 30 minutes. As in the previous usability study, we analyzed participants’ questionnaire responses and reflections to identify high-level trends and opportunities for augmented reading tools.
@@ 5.3. Results
We gathered feedback from participants on various aspects, including the reading experience (immersiveness, understanding, distraction), perceptions of the system (content and query comprehension, trustworthiness, and privacy concerns), and overall usability (content placement and comfort using the MR headset). In general, participants reported feeling more immersed and interested while reading (M=3.5/5, SD=1.04), and they perceived the system as effectively interpreting their queries (M=3.3/5, SD=1.00). Besides, their responses on the content placement within the system were also overall positive (M=3.4/5, SD=1.29).
@@
Participants had varied opinions on whether the system became distracting throughout the reading process (M=2.6/5, SD=1.37) and whether they tended to trust the accuracy and reliability of contents provided by the system (M=2.8/5, SD=1.25). Privacy concerns were also noted by the majority (6/11), and discomfort using the headset was prevalent among the respondents (8/11). These findings are detailed in Figure 7.
Figure 8. Snapshots captured from the Vision Pro during our in-the-wild study.
5.4. Insights
@@
Benefits of Always-On AI Assistant
The exploration of such diverse reading materials can be attributed to the always-on capability of the mixed reality reading assistant, as opposed to screen-based interfaces. A key benefit of this feature is its support for implicit inputs, which dramatically lowers the barrier to asking questions of LLMs. This automatic, continuous content capture broadens the range of applications beyond what participants initially envisioned. For instance, W4 mentioned they had never thought to type a question into ChatGPT when looking at a product label. Even for common reading tasks, they appreciated that the summary was automatically generated and updated in the background as they flipped through pages in a library. Overall, the always-on, implicit assistant made it effortless to integrate AI into daily tasks, expanding potential use cases and drastically reducing the explicit effort required from users.
@@
Participants appreciated the unique affordances offered by mixed reality interfaces, such as the opportunities for spatial and tangible exploration. In a library setting, for example, participants could naturally navigate through bookshelves (spatial exploration), physically pick up a book (tangible interaction), and browse through its pages to automatically receive summaries (always-on assistance). W5 and W10 even leveraged our system to provide summaries or recommendations for entire collections of books on a bookshelf, not just content from a single page. Furthermore, W1 demonstrated the capability to synthesize information across multiple books on a table, requesting summaries or connections between various books and pages. Such observations highlight how mixed reality interfaces enable spatial and tangible interactions that go beyond the limitations of traditional screen-based interfaces, which can only capture and summarize visible content.
@@
We observed several unique and unexpected use cases that took advantage of the LLM’s capacity for free-form questioning and answering. For instance, when reading a physics textbook, W8 asked the system to generate ten quizzes to help him understand quantum mechanics, transforming the textbook into a more active and personalized learning tool. Other common usages included translations (W4, W5, W6), learning aids (W4, W8), and personalized content recommendations (W5, W10). For example, W8 applied the system to his handwritten calculus notes to find answers to particular questions. Beyond traditional reading applications, the AI’s capabilities supported practical decision-making; W4 compared nutritional information of products on store shelves and menu items to determine healthier options, while W5 used it for guidance on waste disposal and printer operations. The LLM’s ability to process a broad spectrum of inquiries clearly expanded the system’s applicability and utility.
@@
Our system offers both 1) proactive assistance through automatic summary generation and 2) on-demand assistance where users manually ask questions. Participants shared their insights on these two functionalities by comparing the benefits and limitations. Proactive summaries were appreciated for providing a useful initial overview of unfamiliar topics. Preferences for proactive summaries varied with reading habits. For example, infrequent readers like W6 and W10 found these proactive summaries helpful in deciding whether to read the content or understand it without reading the full text. On the other hand, W10, who is a more frequent reader, found the constant summary updates unnecessary. In contrast, on-demand question answering allowed for a more focused engagement on specific topics. For instance, W1 enjoyed the “rapid-fire” questioning with a physics textbook, asking more questions than with traditional ChatGPT saying ”You have to describe it properly you have to specifically say what you’re asking about what you’re reading, so that kind of all of this abstracted in this system is like frees up your mind to just focus on the curiosity part Like rapid-fire questions and get rapid-fire answers”. W2 valued how questioning was woven into reading, preserving thought flow without the disruption of toggling between reading and ChatGPT interfaces. Overall, participants favored the on-demand feature due to its applicability to diverse situations, yet many appreciated having both options for different purposes: proactive features for implicit skimming and on-demand for addressing more specific questions.